OpenStudy (anonymous):

Solve and show steps : q + 12 – 2(q – 22) > 0

OpenStudy (anonymous):

q + 12 - 2(q - 22) > 0 q + 12 - 2q - 44 > 0 q -32 - 2q > 0 -32 minus \[2q^{2}\] >0 subract -32 from 0 or add 32 to 0 same thing \[-2q^{2}\] > 32 then square root the 2 from both sides \[\sqrt{-2q}\] > \[\sqrt{32}\] (sorry i cant figure out how to get some of the equations on the same line) once that is done then it will leave u with -2q< \[\sqrt{32}\] you can solve it further but i was tought to leave it it like that if the square root doesnt come out a whole number

OpenStudy (anonymous):

Ok , thanks. :)

OpenStudy (anonymous):

q + 12 - 2q +44 > 0 Distributed the -2 56 - q > 0 grouped like terms q < 56

OpenStudy (anonymous):

@kal94 is incorrect

OpenStudy (anonymous):

you should not have a squared term

OpenStudy (anonymous):

also they did not distribute correctly

OpenStudy (anonymous):

oh ok .

OpenStudy (anonymous):

i was using a calculator and thats what it gave me, ill check to see if i typed in a wrong number or symbol

OpenStudy (anonymous):

q + 12 - 2q +44 > 0 [Distributed the -2] 56 - q > 0 [grouped like terms] q < 56 [added q to each side] To check you can always plug in a number greater than 56 and it shouldn't work and a number less than and it should work.

OpenStudy (anonymous):

@kal94 just look at the equation. the only way to get a q^2 is when there is q*q. This doesn't exist above. Something is wrong

OpenStudy (anonymous):

i realized that now i pressed multiplication symbol instead of plus on the calculator

OpenStudy (anonymous):

That would do it